OpenStudy (anonymous):

write an equation of a line that passes through the points (-2,5) and (1,2)

myininaya (myininaya):

m=(5-2)/(-2-1)=3/(-3)=-1 y=mx+b is a form all lines can be written in y=-1x+b we need to know b we know a point on the line (1,2) x=1 y=2 2=-1(1)+b solve for b 2=2+b b=0 so equation of the line is y=-x

myininaya (myininaya):

oops mistake! 2=-1+b so b=2+1=3 so equation of the line is y=-x+3